OpenStudy (anonymous):

Solve the following inequality and graph the solution set |-4x+9|>14

OpenStudy (anonymous):

-4x+9 > -14 and -4x +9 > 14 -4x> -23 -4x > 5 x < 23/4 x< -5/4

OpenStudy (anonymous):

not to be picky, but it should be OR rather than AND

OpenStudy (anonymous):

actually lets start again

OpenStudy (anonymous):

\[|-4x+9|>14\] \[-4x+9<-14\text{ or } -4x+9>14\] and now we solve these separately

OpenStudy (anonymous):

Aye! I know...I didn't put stress on "and" and "or". But joanne...Satellite is right. It should be or.

OpenStudy (anonymous):

\[-4x+9 <-14\] \[-4x<-23\] \[x>\frac{23}{4}\] or \[-4x+9> 14\] \[-4x>5\] \[x<-\frac{4}{5}\]

OpenStudy (anonymous):

typo on the last one should be \[x<-\frac{5}{4}\]

OpenStudy (anonymous):

so your solution is \[(-\infty, -\frac{5}{4})\cup (\frac{23}{4},\infty)\]

OpenStudy (anonymous):

What would that reduce to cause nothing will go into 23

OpenStudy (anonymous):

5 and 3 what